OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the points (-2, 6) and (-1, 4). y = -2x + 5 y = 2x - 3 y = 2x + 6 y = -2x + 2

OpenStudy (anonymous):

d

OpenStudy (anonymous):

First find the equation of line: y-y1 = (y2-y1)/(x2-x1) x (x-x1) After substituting the given points in the above equation, the equation must be transformed into this form: y = mx +c

OpenStudy (anonymous):

ok thanks